OpenStudy (anonymous):

What is the slope of the line that passes through the points (−5,4) and (15,−4)? A. -2/5 B. 0 C. -5/2 D. undefined

OpenStudy (anonymous):

\[\frac{-4-4}{15-(-5)}=\frac{-8}{20}=-\frac{2}{5}\]

OpenStudy (anonymous):

thanks nancy!